OpenStudy (anonymous):

integral of sin^3 (2x)/ 1+ cos 2x

OpenStudy (thomas5267):

Is this your question? \[ \int\frac{\sin^3(2x)}{1+\cos(2x)}dx \]

OpenStudy (anonymous):

yes

OpenStudy (thomas5267):

Hint: \[ \sin^3(2x)=(\sin(2x))^3\\ \cos(2x)=1-2\sin^2(x) \]

OpenStudy (anonymous):

\[\int\limits(\sin ^{2} 2x)(\sin 2x)=\frac{ \int\limits(1+\cos 2x)(1-\cos2x)(\sin2x }{ 1+\cos2x }\] could this be possible?

hartnn (hartnn):

just wondering an alternative method : s= sin 2x, c = cos 2x s^3 = s*s^2 = s (1-c^2) = s(1+c)(1-c) so we get integral of s(1-c) = integral of s -sc = sin 2x -(1/2) sin 4x

hartnn (hartnn):

aha! pooh, you tried the same thing :) and ofcourse thats possible !

OpenStudy (anonymous):

then is it =-1/2(cos 2x) -1/4(cos 2x) + c

OpenStudy (thomas5267):

I got \(\frac{1}{4}\cos^4(x)+c\)...

hartnn (hartnn):

-1/2(cos 2x) is correct first term, but how did u get -1/4(cos 2x) ?

OpenStudy (anonymous):

im quite confused...can you let u as cos 2x and du= 2 sin 2x? [sin^2 (2x)][sin 2x] / 1+cos 2x dx = [1+cos 2x][1-cos 2x](sin 2x) / 1+cos 2x dx =[1-cos 2x][sin 2x] dx = (sin 2x) - (cos 2x sin 2x) dx =-1/2(cos 2x) -1/4(cos 2x) + c

OpenStudy (anonymous):

du= -2sin 2x

hartnn (hartnn):

i just asked how did u calculate integral (cos 2x sin 2x) dx as 1/4(cos 2x) + c what i would do is (cos 2x sin 2x) dx = 1/2 (2 cos 2x sin 2x) = 1/2 sin 4x dx integrating we get 1/2 (cos 4x)/4 +c or 1/8 cos 4x +c for the 2nd term

hartnn (hartnn):

if you let u as cos 2x and du= 2 sin 2x then (cos 2x sin 2x) dx becomes 1/2 (u) dx integrating u^2/4 +c (cos ^2 2x)/4 +c did you missed the square ?

hartnn (hartnn):

and both are correct and equivalent ways to integrate :)

OpenStudy (thomas5267):

Your way is more clever than mine's.

OpenStudy (thomas5267):

To be honest, I have absolutely no idea what poohtooth is trying to do. Why there is \(\int(\sin^2(2x)\sin(2x))dx\)?

hartnn (hartnn):

\(\sin^3 2x = \sin 2x \times \sin ^2 2x\)

OpenStudy (anonymous):

ah i missed something... i let u=sin2x and du= 2cos2x that's why i got \[\frac{ 1 }{ 2 }\frac{ \cos2x }{ 2 }=\frac{ 1 }{ 4 }\cos 2x\] am i correct?

OpenStudy (anonymous):

@ thomas it is to cancel your denominator..

hartnn (hartnn):

you still get integral u/2 du right ?? integrating it gives you u^2/4 +c ... you doesn't seem to get the square term

OpenStudy (anonymous):

ah ok...so you mean\[(\frac{ \sin 2(2x) }{ 2})\]?

OpenStudy (thomas5267):

I am not sure whether or not you know this but I will just point this out. \[ \int\frac{f(x)}{g(x)}dx{\neq\frac{\int f(x)\,dx}{\int g(x)\,dx}}\neq\frac{1}{g(x)}\int f(x)\,dx \]

OpenStudy (thomas5267):

Which is the integral, a,b or neither? \[ \begin{align*} &\int\frac{\sin^3(2x)}{1+\cos(2x)}dx &\text{a}\\ &\frac{\int\sin^3(2x)\,dx}{1+\cos(2x)}&\text{b} \end{align*} \]

OpenStudy (anonymous):

a

OpenStudy (thomas5267):

Have you learned integration by substitution?

OpenStudy (anonymous):

yes of course...